a) The man pulled a suitcase 190 meters along the airport floor. The magnitude of the tensile force was 22 N and formed an angle of 55° with the horizontal. How much work did the Driving Force do?


b) At the same time, the man was holding a mobile phone weighing 143 g in his ear. How much work did the power needed to support a cell phone do?


1
Expert's answer
2021-01-19T13:20:16-0500

a) By the definition of the work done, we have:


"W=Fdcos\theta,""W=22\ N\cdot190\ m\cdot cos55^{\circ}=2397.5\ J."


b) Since the only one force acts on the mobile phone (the force of gravity), and mobile phone doesn't move in vertical direction, the work required to support the mobile phone equals zero.

Answer:

a) "W=22\ N\cdot190\ m\cdot cos55^{\circ}=2397.5\ J."

b) "W=0".